OpenStudy (mahogany):

A 8.6 kg block initially at rest is pulled to the right along a horizontal, frictionless surface by a constant, horizontal force of 16.3 N. Find the speed of the block after it has moved 4.3 m. Answer in units of m/s.

OpenStudy (irishboy123):

Equate Work done with (kinetic) energy gained. \(W = F*x = \frac{1}{2} m v^2\)